the length and width of a rectangle are in the ratio of 3 to 3.if the perimeter is 80cm,calculate.the width,the length of a diagonal, the area of the rectangle
Answer by opeodu(24)

L/b=3/3 l=b 2(l+b)=80 l+b=40 2b=40 b=20cm l=20cm length of diagonal=sqrt 20^2+20^2=sqrt800 area= 20*20=400cm2